OpenStudy (anonymous):

Look at the following points. (4,0)(3,-1)(6,3)(2,-4) Which are solutions to y=x-4? choose all correct answers. ( 2 points ) 6,3 4,0 3,-1 2,-4

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

I got (4,0) y = 4 - 4 y = 0

OpenStudy (anonymous):

I also got (3,-1) y = 3 - 4 y = -1